OpenStudy (anonymous):

Four trumpet players’ instruments are mixed up, and the trumpets are given to the players just before a concert. What is the probability that no one gets his or her trumpet back?

OpenStudy (anonymous):

I know that the answer is 3/8 but I don't know why. HELP!

OpenStudy (anonymous):

are you sure it's 3/8?

OpenStudy (anonymous):

total number of ways to distribute = 4!=4*3*2=24 no one gets the correct one = 3+3+3 (i.e., each one can have 3 other ones) P=(3+3+3)/24=3*3/24=3/8